# MigGPT: Harnessing Large Language Models for Automated Migration of Out-of-Tree Linux Kernel Patches Across Versions

### v2

@@ Abstract
Out-of-tree kernel patches are essential for adapting the Linux kernel to new hardware or enabling specific functionalities. Maintaining and updating these patches across different kernel versions demands significant effort from experienced engineers. Large language models (LLMs) have shown remarkable progress across various domains, suggesting their potential for automating out-of-tree kernel patch migration. However, our findings reveal that LLMs, while promising, struggle with incomplete code context understanding and inaccurate migration point identification. In this work, we propose MigGPT, a framework that employs a novel code fingerprint structure to retain code snippet information and incorporates three meticulously designed modules to improve the migration accuracy and efficiency of out-of-tree kernel patches. Furthermore, we establish a robust benchmark using real-world out-of-tree kernel patch projects to evaluate LLM capabilities. Evaluations show that MigGPT significantly outperforms the direct application of vanilla LLMs, achieving an average completion rate of 74.07% ($\uparrow 45.92\%$) for migration tasks.
@@ 1 Introduction
Existing code migration technologies diff_catchup ; a3 ; app_evovle ; cocci_evolve ; neat ; T2R ; REFAZER ; PYEVOLVE ; SKYPORT ; FixMorph ; PatchWeave ; TSBPORT utilize static program analysis static_program_analysis to facilitate API cross-version maintenance or the backporting of CVE security patches. However, these methods only address a subset of scenarios in out-of-tree kernel patch migration. 1) They rely on predefined migration rules, which are insufficient for handling comprehensive scenarios involving complex changes such as namespace modifications, invocation conflict resolution, and the integration of control and data flow dependencies.
@@
2) These single-step methods assume known target code locations in updated repositories, limiting their applicability to out-of-tree kernel patch scenarios.
@@ 2.2 LLMs for Coding
3 Problem Formulation
[…]
4 Migration Benchmark
4.1 Migration Types
@@ 5.3 Code Fingerprint
To address the limitations of traditional code structure analysis, we propose Code Fingerprint (CFP), a lightweight sequential data structure for analyzing code snippets. CFP records both the content and positional information for each line of statements, encompassing all C language statements, including comments and inline assembly (a detailed example is provided in App. D.1). As shown in Figure 4, CFP focuses on recording function definitions and function calls, which are crucial for addressing challenges 1 and 3, as detailed in Section 5.4. Additionally, its linear structure facilitates accurate positioning for insertion, deletion, and other update operations, tackling challenges 2 and 4, further explained in Sections 5.5 and  5.6. The algorithm for generating CFP is in App. D.4. Overall, CFP offers three key advantages: 1) effective processing of incomplete code snippets, 2) preservation of critical information such as comments and inline assembly, which are vital for out-of-tree kernel patch migration, and 3) a streamlined design that focuses on essential statements, improving migration efficiency and reducing overhead. CFP is specifically designed for out-of-tree kernel patch code, addressing the challenges encountered in Linux kernel patches. More discussions about CFP are in App. G.2. By minimizing unnecessary processing while ensuring relevance, CFP provides a targeted solution for migrating out-of-tree kernel patches across Linux kernel versions.
@@ 5.4 Retrieval Augmentation Module
The retrieval augmentation module is designed to address challenge 1 and challenge 3 encountered during the migration update of out-of-tree kernel patches by LLMs. In challenge 1, LLMs are prone to be misled by similar function structure when processing function definitions in code snippets, leading to incorrect retrieval of $s_{\text{new}}$ in $\text{file}_{\text{new}}$, which ultimately results in erroneous migrated $s^{\prime}_{\text{new}}$. To overcome this challenge, it is necessary to constrain the LLM’s attention to the target code snippet. As illustrated in Figure 2, the retrieval augmentation module achieves this by constructing a code fingerprint structure ($\text{CFP}_{\text{old}}$) for the old version of the Linux kernel code snippets $s_{\text{old}}$. By analyzing $\text{CFP}_{\text{old}}$, the module extracts the function signatures of the function definitions contained within $s_{\text{old}}$. These function signatures are then used to build a prompt to describe the structure information (“Structure Prompt"), which is incorporated into the input fed to the LLM. An example is provided in the App. D.2 Figure 8.
@@ 6.1 Evaluation Settings
We assess MigGPT using two benchmarks: the out-of-tree kernel patch migration benchmark from Section 4 and FixMorph’s CVE patch backporting benchmark FixMorph , which includes 350 instances. For baselines, we use vanilla LLMs, including GPT-3.5 gpt-3.5 , GPT-4-turbo gpt4 , OpenAI-o1 openai-o1 , DeepSeek-V2.5 deepseek-2.5 , DeepSeek-V3 deepseek-3 , Deepseek-R1 deepseek-r1 , Llama-3.1-8B and Llama-3.1-70B-Instruct llama-3.1 , as they are widely recognized for their advanced performance, along with previous migration efforts like FixMorph FixMorph , TSBPORT TSBPORT and PPatHF PPatHF . Evaluation metrics include “best match" (exact code similarity after removing spaces, line breaks, and tab characters), “semantic match" (CodeBLEU with a 0.9 threshold for binary classification, detailed in App. F.5codebleu , and “human match" (developer-judged functional equivalence, detailed in App. F.2). We also considered the compilation success rate, as detailed in App. F.3. The hyperparameter $m$ is set to 3.
@@
Table 2: The accuracy of the MigGPT-augmented LLMs compared to vanilla LLMs in retrieving target code snippets.
@@
Table 3: The accuracy of the MigGPT-augmented LLMs compared to vanilla LLMs in the migration task of target code snippets.
@@ 6.2 Performance Evaluation (RQ1)
MigGPT outperformed previous migration efforts.
We also evaluate the performance of MigGPT in the context of CVE patch backporting. Notably, our out-of-tree patch migration task is different from FixMorph, TSBPORT, and PPatHF, which only solve the migration problem, while we target both target code retrieval and migration. As illustrated in Table 5, MigGPT demonstrates superior performance compared to previous patch backporting methods. We also discuss the generalization of MigGPT in App. G.3.
@@ 6.3 Ablation Study (RQ2)
We conduct an ablation study to evaluate the impact of the four units in MigGPT: CFP, Retrieval Augmentation Module, Retrieval Alignment Module, and Migration Augmentation Module (details in App. F.6). Figure 6 presents the outcomes of four tested variants on our benchmarks. Among these, MigGPT consistently outperforms the ablation baselines. Meanwhile, we perform an ablation study on hyperparameter $m$ in Algorithm 1 with MigGPT-GPT-4-turbo, which controls the total query time of the Retrieval Augmentation Module. As shown in Figure 6, $m=3$ is suitable for both Type 1 and Type 2 examples.
@@ 6.4 Failure Analysis (RQ3)
Table 7: The number of line edit distances between failed cases and human-migrated outcomes in the migration task. $3\leq\text{dis}<6$” denotes a line edit distance of at least 3 but less than 6.
@@
We evaluate MigGPT’s robustness by analyzing failed migration cases (not human-matched) across various samples, measuring line edit distances (insertions, deletions, modifications) between MigGPT’s incorrect outputs and human-corrected results (see App. F.4). As shown in Table 7, 41% of MigGPT’s errors require fewer than three lines of modification to align with correct results, demonstrating its potential to aid in out-of-tree kernel patch migration.
@@ A.2 Examples of Benchmark
Class Type 1 Type 2 Formalization $\Delta\neq\varnothing,\Sigma\neq\varnothing,$ $\Delta\neq\varnothing,\Sigma\neq\varnothing,$ $\forall\delta\in\Delta,\forall\sigma\in\Sigma,\langle\delta,\sigma\rangle\neq 0$ Number 80 (48.2%) 86 (51.8%) $s_{\text{old}}$ vs $s^{\prime}_{\text{old}}$ [Uncaptioned image] [Uncaptioned image] $s_{\text{old}}$ vs $s_{\text{new}}$ [Uncaptioned image] [Uncaptioned image]
@@ D.4 Algorithm and Prompts
Algorithm 3 illustrates the specific details about generating CFP. The code snippet $s$ is tokenized, and nested scopes (e.g., {,}, #ifdef/#endif) are identified via bracket matching. Critical symbols (e.g., {, }, #ifdef, func()) demarcate code blocks. Function calls (e.g., foo()) are detected through pattern matching on token sequences (e.g., identifier followed by (). Associated functions are identified by analyzing call statements within code blocks, avoiding call graph construction.
@@ Appendix E More Results of MigGPT
Table 9: The accuracy of the MigGPT-augmented LLMs compared to vanilla LLMs in retrieving target code snippets.
@@
Table 10: The accuracy of the MigGPT-augmented LLMs compared to vanilla LLMs in the migration task of target code snippets.
@@ F.1 Platform
F.2 The Reliability of Human Match
@@
F.3 Compilation Success Rate
@@
F.4 Line Edit Distance
@@ G.2 CFP and Intermediate Representations
  • Context Preservation: Kernel patches often contain conditional compilation macros, inline assembly, and annotations stripped during IR generation (e.g., preprocessing eliminates macros). CFP retains these alongside code structure (function signatures, control-flow anchors) to provide LLMs with full migration context.
@@ G.3 Generalization of MigGPT
G.4 Impact Statement

### v3

@@ Abstract
Out-of-tree kernel patches are essential for adapting the Linux kernel to new hardware or enabling specific functionalities. Maintaining and updating these patches across different kernel versions demands significant effort from experienced engineers. Large language models (LLMs) have shown remarkable progress across various domains, suggesting their potential for automating out-of-tree kernel patch migration. However, our findings reveal that LLMs, while promising, struggle with incomplete code context understanding and inaccurate migration point identification. In this work, we propose MigGPT, a framework that employs a novel code fingerprint structure to retain code snippet information and incorporates three meticulously designed modules to improve the migration accuracy and efficiency of out-of-tree kernel patches. Furthermore, we establish a robust benchmark using real-world out-of-tree kernel patch projects to evaluate LLM capabilities. Evaluations show that MigGPT significantly outperforms the direct application of vanilla LLMs, achieving an average completion rate of 74.07% ($\uparrow 45.92\%$) for migration tasks. Our code and data are available at https://github.com/CherryBlueberry/MigGPT.
@@ 1 Introduction
Existing code migration technologies diff_catchup ; a3 ; app_evovle ; cocci_evolve ; neat ; T2R ; REFAZER ; PYEVOLVE ; SKYPORT ; FixMorph ; PatchWeave ; TSBPORT utilize static program analysis static_program_analysis to facilitate API cross-version maintenance or the backporting of CVE security patches. However, these methods only address a subset of scenarios in out-of-tree kernel patch migration. 1) They rely on predefined migration rules, which are insufficient for handling comprehensive scenarios involving complex changes such as namespace modifications, invocation conflict resolution, and the integration of control and data flow dependencies. 2) These single-step methods assume known target code locations in updated repositories, limiting their applicability to out-of-tree kernel patch scenarios.
@@ 2.2 LLMs for Coding
3 Problem Formulation
[…]
4 Migration Benchmark
4.1 Migration Types
@@ 5.3 Code Fingerprint
To address the limitations of traditional code structure analysis, we propose Code Fingerprint (CFP), a lightweight sequential data structure for analyzing code snippets. CFP records both the content and positional information for each line of statements, encompassing all C language statements, including comments and inline assembly (a detailed example is provided in App. D.5). As shown in Figure 4, CFP focuses on recording function definitions and function calls, which are crucial for addressing challenges 1 and 3, as detailed in Section 5.4. Additionally, its linear structure facilitates accurate positioning for insertion, deletion, and other update operations, tackling challenges 2 and 4, further explained in Sections 5.5 and  5.6. The algorithm for generating CFP is in App. D.5. Overall, CFP offers three key advantages: 1) effective processing of incomplete code snippets, 2) preservation of critical information such as comments and inline assembly, which are vital for out-of-tree kernel patch migration, and 3) a streamlined design that focuses on essential statements, improving migration efficiency and reducing overhead. CFP is specifically designed for out-of-tree kernel patch code, addressing the challenges encountered in Linux kernel patches. More discussions about CFP are in App. G.2. By minimizing unnecessary processing while ensuring relevance, CFP provides a targeted solution for migrating out-of-tree kernel patches across Linux kernel versions.
@@ 5.4 Retrieval Augmentation Module
The retrieval augmentation module is designed to address challenge 1 and challenge 3 encountered during the migration update of out-of-tree kernel patches by LLMs. In challenge 1, LLMs are prone to be misled by similar function structure when processing function definitions in code snippets, leading to incorrect retrieval of $s_{\text{new}}$ in $\text{file}_{\text{new}}$, which ultimately results in erroneous migrated $s^{\prime}_{\text{new}}$. To overcome this challenge, it is necessary to constrain the LLM’s attention to the target code snippet. As illustrated in Figure 2, the retrieval augmentation module achieves this by constructing a code fingerprint structure ($\text{CFP}_{\text{old}}$) for the old version of the Linux kernel code snippets $s_{\text{old}}$. By analyzing $\text{CFP}_{\text{old}}$, the module extracts the function signatures of the function definitions contained within $s_{\text{old}}$. These function signatures are then used to build a prompt to describe the structure information (“Structure Prompt"), which is incorporated into the input fed to the LLM. An example is provided in the App. D.2.
@@ 6.1 Evaluation Settings
We assess MigGPT using two benchmarks: the out-of-tree kernel patch migration benchmark from Section 4 and FixMorph’s CVE patch backporting benchmark FixMorph , which includes 350 instances. For baselines, we use vanilla LLMs, including GPT-3.5 gpt-3.5 , GPT-4-turbo gpt4 , OpenAI-o1 openai-o1 , DeepSeek-V2.5 deepseek-2.5 , DeepSeek-V3 deepseek-3 , Deepseek-R1 deepseek-r1 , Llama-3.1-8B and Llama-3.1-70B-Instruct llama-3.1 , as they are widely recognized for their advanced performance, along with previous migration efforts like FixMorph FixMorph , TSBPORT TSBPORT and PPatHF PPatHF . Evaluation metrics include “best match" (exact code similarity after removing spaces, line breaks, and tab characters), “semantic match" (CodeBLEU with a 0.9 threshold for binary classification, detailed in App. F.6codebleu , and “human match" (developer-judged functional equivalence, detailed in App. F.3). We also considered the compilation success rate, as detailed in App. F.4. The hyperparameter $m$ is set to 3.
@@
Table 2: The accuracy of the MigGPT-augmented LLMs compared to vanilla LLMs on retrieving target code snippets.
@@
Table 3: The accuracy of the MigGPT-augmented LLMs compared to vanilla LLMs on the migration task of target code snippets.
@@ 6.2 Performance Evaluation (RQ1)
MigGPT outperformed previous migration efforts on the CVE patch backporting task.
We also evaluate the performance of MigGPT in the context of CVE patch backporting. Notably, our out-of-tree patch migration task is different from FixMorph, TSBPORT, and PPatHF, which only solve the migration problem, while we target both target code retrieval and migration. As illustrated in Table 5, MigGPT demonstrates superior performance compared to previous patch backporting methods.
Table 6: The accuracy of MigGPT compared to vanilla LLM on the Linux driver migration task.
Method GPT-4-turbo DeepSeek-V3 Vanilla MigGPT Vanilla MigGPT Best Match 54.16% 83.33% (+29.17%) 58.33% 79.16% (+20.83%) Semantic Match 62.50% 83.33% (+20.83%) 62.50% 87.50% (+25.00%)
MigGPT performs well on the Linux driver migration task.
We conducted experiments to assess MigGPT’s performance on the Linux driver migration. We randomly collected 24 driver migration samples. The results, presented in the Table 6, indicate that MigGPT performs well in migrating patches involving driver interface modifications.
Table 7: The accuracy of MigGPT compared to vanilla LLM on Python code migration task.
Method GPT-4-turbo DeepSeek-V3 Vanilla MigGPT Vanilla MigGPT Best Match 61.29% 80.65% (+19.36%) 64.52% 83.87% (+19.35%) Semantic Match 70.97% 87.10% (+16.13%) 74.19% 90.32% (+16.13%)
MigGPT demonstrates cross-language generalizability.
We conducted an experiment focusing on Python code migration. We randomly selected 31 migration examples from our benchmark and translated them to Python. With only minor adjustments (modify the implementation of the statement tokenization and statement element extraction functions to adapt the CFP for Python syntax) MigGPT can process these Python examples. The results of Table 7 demonstrate that the average semantic match performance improved by 16.13%. compared to a vanilla approach, indicating MigGPT’s effectiveness beyond C-language code. We also discuss the generalization of MigGPT in App. G.3.
@@ 6.3 Ablation Study (RQ2)
6.4 Ablation Study
We conduct an ablation study to evaluate the impact of the four units in MigGPT: CFP, Retrieval Augmentation Module, Retrieval Alignment Module, and Migration Augmentation Module (details in App. F.7). Figure 6 presents the outcomes of four tested variants on our benchmarks. Among these, MigGPT consistently outperforms the ablation baselines. Meanwhile, we perform an ablation study on the hyperparameter $m$ in Algorithm 1 with MigGPT-GPT-4-turbo, which controls the total query time of the Retrieval Augmentation Module. As shown in Figure 6, $m=3$ is suitable for both Type 1 and Type 2 examples.
@@ 6.5 Failure Analysis (RQ3)
Table 9: The line edit distance between the failure cases of MigGPT-augmented GPT-4-Turbo and the manual migration results. $3\leq\text{dis}<6$” denotes a line edit distance of at least 3 but less than 6.
@@
We evaluate MigGPT’s robustness by analyzing failed migration cases (not human-matched) across various samples, measuring line edit distances (insertions, deletions, modifications) between MigGPT’s incorrect outputs and human-corrected results (see App. F.5). As shown in Table 9, 41% of MigGPT’s errors require fewer than three lines of modification to align with correct results, demonstrating its potential to aid in out-of-tree kernel patch migration.
In our in-depth analysis of "human match" errors in MigGPT-augmented GPT-4-turbo, we identified the following primary categories:
  • Incomplete $s_{new}$ Retrieval for Large Codebases (31.03%): In some instances, when retrieving $s_{new}$ from $\text{file}_{new}$, the target code $s_{new}$ was significantly large (exceeding 150 lines), leading to incomplete retrieval. This could be due to LLMs’ tendency to shift attention when dealing with long contexts.
  • Deviation from the Migration Point in $s^{\prime}_{new}$ Generation (24.14%): Even with precise migration point information provided by CFP, the LLM occasionally failed to strictly adhere to these locations during $s^{\prime}_{new}$ generation. While minor offsets often didn’t impact functionality, they sometimes led to functional errors in code with complex control or data flows. This behavior appears to be related to the inherent randomness of LLMs.
  • Difficulty in Fusing Divergent Changes in Type 2 Migrations (27.59%): For certain Type 2 migrations, significant differences between $s_{new}$ and $s_{old}$’s modifications to $s_{old}$ prevented the LLM from correctly integrating these changes, resulting in errors. This limitation points to challenges related to the LLM’s code comprehension and manipulation capabilities.
  • Miscellaneous (17.24%): Errors in symbols, statements, etc., appearing in $s^{\prime}_{new}$, such as incorrectly writing verbose(env, off, size, reg’s id) instead of verbose(env, off, size, reg->id). This may be related to LLM hallucinations.
It’s important to note that these identified failure modes primarily arise from the inherent limitations of the LLMs themselves, rather than architectural flaws within the MigGPT framework.
@@ 7 Conclusion
Acknowledgments and Disclosure of Funding
This work is partially supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (Grants No.XDB0660300, XDB0660301, XDB0660302), Science and Technology Major Special Program of Jiangsu (Grant No. BG2024028), the NSF of China (Grants No. U22A2028, 62302483, 6240073476), CAS Project for Young Scientists in Basic Research (YSBR-029) and Youth Innovation Promotion Association CAS.
@@ A.2 Examples of Benchmark
Class Type 1 Type 2 Formalization $\Delta\neq\varnothing,\Sigma\neq\varnothing,$ $\Delta\neq\varnothing,\Sigma\neq\varnothing,$ $\forall\delta\in\Delta,\forall\sigma\in\Sigma,\langle\delta,\sigma\rangle\neq 0$ Number 80 (59.3%) 55 (40.7%) $s_{\text{old}}$ vs $s^{\prime}_{\text{old}}$ [Uncaptioned image] [Uncaptioned image] $s_{\text{old}}$ vs $s_{\text{new}}$ [Uncaptioned image] [Uncaptioned image]
@@ D.4 Algorithm and Prompts
D.5 CFP
Algorithm 3 illustrates the specific details about generating CFP. The code snippet $s$ is tokenized, and nested scopes (e.g., {,}, #ifdef/#endif) are identified via bracket matching. Critical symbols (e.g., {, }, #ifdef, func()) demarcate code blocks. Function calls (e.g., foo()) are detected through pattern matching on token sequences (e.g., identifier followed by (). Associated functions are identified by analyzing call statements within code blocks, avoiding call graph construction. An example of step-by-step CFP generation is illustrated in Figure 16.
@@ Appendix E More Results of MigGPT
Table 11: The accuracy of the MigGPT-augmented LLMs compared to vanilla LLMs on retrieving target code snippets.
@@
Table 12: The accuracy of the MigGPT-augmented LLMs compared to vanilla LLMs on the migration task of target code snippets.
@@ F.1 Platform
F.2 Contextual Information
The contextual information given to the LLM is identical for both the “two-step strategy” (baseline vanilla approach in Tables 1 and 2) and the “one-step strategy”, so the comparison is fair (details are shown in Table 13). The contextual information includes the old version code $s_{old}$, $s\prime_{old}$, and the new version file new.c. As shown in Figure 5 of the paper, the “two-step strategy” is clearly superior to the “one-step strategy”.
Table 13: The contextual information of various methods.
Context One-step strategy Two-step strategy (Vanilla) One-step strategy + CFP MigGPT (Two-step strategy + CFP) Code Information CFP Information
We provided the key migration information extracted by CFP and the code information ($s_{old}$, $s\prime_{old}$, new.c) as context to the “one-step strategy” (One-step + CFP) and compared it with MigGPT (Two-step + CFP). As shown in Table 14, formulating the migration task as a two-step process demonstrates performance advantages.
Table 14: The accuracy of MigGPT and “One-step + CFP” on the migration task.
Method Best Match Semantic Match One-step + CFP 38.52% 42.96% MigGPT 62.96% 80.00%
F.3 The Reliability of Human Match
@@
F.4 Compilation Success Rate
@@
F.5 Line Edit Distance
@@ G.2 CFP and Intermediate Representations
  • Context Preservation: Kernel patches often contain conditional compilation macros, inline assembly, and annotations stripped during IR generation (e.g., preprocessing eliminates macros). CFP retains these alongside code structure (function signatures, control-flow anchors) to provide LLMs with a full migration context.
@@ G.3 Generalization of MigGPT
G.4 Structured Analysis of CFP
The high-level idea of using program analysis techniques to create structured representations of code to guide LLMs is an emerging area of research. We have compared MigGPT with existing research in this emerging area. For instance, works such as AutoOS [3] and BYOS [29] utilize heuristic tree structures to assist LLMs in optimizing kernel configurations during the deployment of operating systems. Similarly, LLM-based fuzzing tools [49, 18] employ formal templates to guide LLMs in generating more effective test cases. These approaches, much like MigGPT, leverage deterministic information such as trees, graphs, and lists to mitigate the inherent uncertainty and randomness of LLMs, thereby improving performance across various application tasks. However, it’s worth noting that the specific deterministic information used and its data structure are often tailored to the particular task at hand.
G.5 Deprecated System Calls
Once a system call is added to the kernel, it’s generally supported indefinitely to avoid breaking existing software that relies on it. However, kernels do evolve, introducing new system calls that offer more robust, secure, or efficient functionalities. These new features can, in some cases, functionally supersede older, more limited mechanisms. When a patch migration involves new system calls introduced in a newer kernel version, MigGPT is well-equipped to handle such scenarios. In these cases, the new version target code ($s_{new}$) will contain examples of how the new system call is used. MigGPT can then refer to these usage examples in $s_{new}$ to complete the migration and generate a corresponding patch that aligns with the new API. Certainly, in this scenario, migration types can still be categorized according to the rules in Table 1. If the new kernel version and patch modifications don’t overlap in location, it’s Type 1. Otherwise, it’s Type 2.
G.6 Impact Statement